OpenStudy (anonymous):

Prove that (sectheta)(csctheta)(cottheta)=csc^2theta

OpenStudy (australopithecus):

csc = 1/sin sec = 1/cos cot = 1/tan

OpenStudy (australopithecus):

tan = sin/cos thus cot = cos/sin

OpenStudy (australopithecus):

using these proofs you can easily break this up and solve it using basic algebra

OpenStudy (callisto):

\[LHS = sec\theta csc\theta cot \theta = sec\theta csc\theta \frac{cos\theta}{sin\theta}\]\[ = sec\theta csc\theta cos\theta csc\theta = csc^2\theta = RHS\]

OpenStudy (campbell_st):

remember cot = cos/sin \[\frac{1}{\cos(\theta)}\times \frac{1}{\sin(\theta)} \times \frac{\cos(\theta)}{\sin(\theta)}\] this its solution is \[\frac{1}{\sin(\theta)\times \sin(\theta)} = \frac{1}{sin^2(theta)}=csc^2(theta)\]